# Enhancing Morphological Measurements of Cosmic Web with Delaunay Tessellation Field Estimation

### v1

@@ 1 Introduction
Ambitious on-going and up-coming cosmological surveys [e.g., HETDEX (Hill et al. 2008), LSST (LSST Science Collaboration et al. 2009), Euclid (Laureijs et al. 2011), 4MOST (de Jong et al. 2012), PFS (Takada et al. 2014), SPHEREX (Doré et al. 2014), DESI (DESI Collaboration et al. 2016), WFIRST (Doré et al. 2018), and CSST (Gong et al. 2019)], particularly the fifth-generation surveys [e.g., WST (Ellis & Dawson 2019), MSE (Percival et al. 2019), MegaMapper (Schlegel et al. 2019), and MUST [^1] MUltiplexed Survey Telescope (MUST): https://must.astro.tsinghua.edu.cn/must/en/index.html], will provide high-precision map of cosmic large-scale structure (LSS), which encodes a wealth of valuable cosmological information about our Universe. Efficient extractions of these critical information will help us greatly deepen our understanding of many key fundamental questions in cosmology, (e.g., dark energy properties, neutrino masses, gravities and inflation models, etc.). This necessitates the development of powerful statistical tools to characterize or quantify LSS properties from various angles.
@@
Consequently, various summary statistics for non-Gaussian information have been proposed as potential supplements to two-point statistics, e.g., count in cells (de Lapparent et al. 1991; Uhlemann et al. 2020; Repp & Szapudi 2020), void statistics (Chan et al. 2014; Zhao et al. 2016; Kreisch et al. 2019), peak statistics (Gay et al. 2012; Shan et al. 2014; Liu et al. 2015a; Liu et al. 2015b), Voronoi statistics (Paranjape & Alam 2020; Jamieson & Loverde 2021), and even scattering transform (Cheng et al. 2020; Cheng & Ménard 2021), etc. In particular, morphological statistical methods [e.g., $\beta$-Skeleton (Fang et al. 2019; Suárez-Pérez et al. 2021), Betti numbers (Pranav et al. 2017; Pranav et al. 2019; Giri & Mellema 2021), persistent topology (Elbers & van de Weygaert 2019, 2023; Wilding et al. 2021; Jalali Kanafi et al. 2023; Bermejo et al. 2024; Yip et al. 2024), minimal spanning tree (Barrow et al. 1985; Naidoo et al. 2021), cosmic web skeleton (Novikov et al. 2006; Sousbie et al. 2008), wavelet analyses (Martinez et al. 1993; Arnalte-Mur et al. 2012), shape statistics (Sahni et al. 1998; Basilakos 2003), genus statistics (Gott et al. 1986; Hamilton et al. 1986; Appleby et al. 2021a), Minkowski tensors (Appleby et al. 2018a; Appleby et al. 2018b), Minkowski functionals (Mecke et al. 1994; Liu et al. 2020; Matsubara et al. 2022), etc.] are actively employed to characterize the geometrical and topological properties of LSS, providing alternative ways to capture higher-order information directly, in complementary to traditional $N$-point formalism.
@@
As a conceptual generalization of genus (Mecke et al. 1994; Schmalzing & Buchert 1997), Minkowski functionals (MFs) can elegantly describe the global characterizations of morphological properties of LSS, and have been employed in various cosmological studies [e.g., detecting primordial non-Gaussianities (Hikage et al. 2006; Hikage et al. 2008; Planck Collaboration et al. 2016; Ducout et al. 2013), serving as standard ruler (Park & Kim 2010; Zunckel et al. 2011; Blake et al. 2014), constraining cosmological parameters (Appleby et al. 2018c; Appleby et al. 2021a; Appleby et al. 2021b), probing neutrino masses (Marques et al. 2019; Liu et al. 2020; Liu et al. 2022), analyzing effects of redshift space distortions (Codis et al. 2013; Jiang et al. 2021), testing cosmologies and gravities (Hikage et al. 2003; Wang et al. 2012; Fang et al. 2017), etc.]. One popular method [other method can also be found in literatures, i.e., the germ-grain approach (cf. Mecke et al. 1994; Wiegand et al. 2014)] for measuring MFs is the iso-density approach, which we will focus on in our study (see Section 2). The MFs measured in this way can be well theoretically modelled for Gaussian (Tomita 1990; Schmalzing & Buchert 1997; Matsubara 2003) and weakly non-Gaussian (Matsubara 1994; Matsubara 2003; Pogosyan et al. 2009; Matsubara 2010; Gay et al. 2012; Matsubara & Kuriki 2021; Matsubara et al. 2022) fields [even for the fields with RSDs (Matsubara 1996; Codis et al. 2013)], showing distinct advantages over the germ-grain approach.
@@
In this study, we opt for Delaunay Tessellation Field Estimator (DTFE) [^3] Delaunay Tessellation Field Estimator (DTFE): https://github.com/MariusCautun/DTFE (Schaap & van de Weygaert 2000; van de Weygaert & Schaap 2009; van de Weygaert et al. 2009; Cautun & van de Weygaert 2011a; Cautun & van de Weygaert 2011b) to address this issue [^4] Previously, Martínez et al. 2005 attempted to use wavelet-denoising method to alleviate the problem induced by fixed-kernel smoothing. Additionally, Zhang et al. 2010 endeavored to boost topological information of genus statistics with DTFE.. This method can help to recover the largest number of structural elements from point sets, allowing for extracting maximum amount of morphological information with MFs. Therefore, the statistical power of MFs with DTFE (hereafter DTFE MFs) can be significantly enhanced, over the fixed smoothing methods (Zhang et al. 2010). We first demonstrate this at matter field level and further extend it to halo field level. In particular, for DTFE MFs of halo fields, we numerically explore various main systematic effects, caused by finite voxel size, halo number density, halo weighting scheme, and RSDs, on their measurements, and showcase their strong discriminative and constraining power in RSD studies. Hopefully, the applications of DTFE will dramatically improve the performance of MFs in extracting various key cosmological information (e.g., neutrino masses (Liu et al. 2020), modified gravities (Fang et al. 2017), primordial non-Gaussianities (Hikage et al. 2006; Hikage et al. 2008), and cosmological parameters (Appleby et al. 2020; Appleby et al. 2022), etc.) from sparse tracers.
@@ 2 Minkowski Functionals
MFs are a family of morphological (i.e., geometrical and topological) descriptors with properties of additivity, motion invariance, and conditional continuity for any manifold in $D$-dimensional space. These descriptors are originally derived from the theory of convex bodies and integral geometry, and were first introduced into cosmology by Ref. Mecke et al. 1994 to characterize the morphology (i.e., geometry and topology) of cosmic web. According to Hadwiger’s theorem (Hadwiger Springer, Berlin, 1957), the topo-geometrical properties of any given manifold in $D$-dimensional space can be completely characterized by $D+1$ MFs. Thus it opens a unique way to comprehensively access all orders of N-point correlation information at once.
@@
Two standard grid-based numerical algorithms to compute MFs from regularly gridded density field [^5] In literatures, several novel triangulation-based algorithms (cf. Sheth et al. 2003; Aragon-Calvo et al. 2010; Lippich & Sánchez 2021), which estimate MFs by using triangulated iso-density surfaces, also have been proposed in succession, aiming to improve the accuracies of MFs’ estimations., i.e., Koenderink invariant from differential geometry (Schneider 1993) and Crofton’s formula from integral geometry (Hadwiger Springer, Berlin, 1957; Mecke 1979), have been developed in literatures (Schmalzing & Buchert 1997). In this work, we employ Crofton’s formula to quote our results, because it is more stable and is the most commonly used method in previous works. All MFs are measured as functions of density threshold $\nu\equiv 1+\delta$ with $\nu\in\left[0.003,1000\right]$, where $\delta$ is the density contrast. The error bars are estimated via the standard errors of MFs, i.e., $s_{\mathrm{e}}=\sigma/\sqrt{n_{\mathrm{f}}}$. Here, $\sigma$ is the standard deviation of MFs, calculated from $n_{\mathrm{f}}$ subfields obtained by equally subdividing the original field. For DM and halo fields, the numbers of subfields are $n_{\mathrm{f}}=4^{3}=64$ and $n_{\mathrm{f}}=8^{3}=512$, respectively. Since the error bars are too tiny to be visible in most scenarios, they are omitted in all figures in this paper, except for Fig. 12 and Fig. 13. Note that all MFs are visualized by logarithmic x-axis, considering that probability distribution function of LSS roughly obeys log-normal form at low redshift.
@@ 4 Cosmological $N$-body Simulations and Data Samples
In this work, we adopt two high-resolution pure cold dark matter (DM) $N$-body simulations for different application requirements. One simulation is dubbed as WMAP_$\_3072\_600$, which is from the CosmicGrowth simulation suite (Jing 2019) realized by running an adaptive parallel particle–particle–particle-mesh ($\mathrm{P}^{3}\mathrm{M}$) $N$-body code (Jing & Suto 2002; Jing et al. 2007). This simulation incorporates $N_{\mathrm{p}}=3072^{3}$ DM particles with mass resolution of $5.5\times 10^{8}\,h^{-1}M_{\odot}$ in a periodic cubic box with size of $L=600\,h^{-1}\mathrm{Mpc}$, adopting a WMAP cosmology, i.e., $[\Omega_{\mathrm{c}}$, $\Omega_{\mathrm{b}}$, $h$, $n_{\mathrm{s}}$, $\sigma_{8}]$ = [0.2235[0.2235[ 0.2235, $0.0445$, $0.71$, $0.968$, 0.83]0.83]0.83 ]. Another one is called TianZero simulation (Emberson et al. 2017; Yu et al. 2017), realized by using a publicly-available $\mathrm{P}^{3}\mathrm{M}$ $N$-body code CUBEP3M (Harnois-Déraps et al. 2013). This simulation, parameterized with $[\Omega_{\mathrm{c}}$, $\Omega_{\mathrm{b}}$, $h$, $n_{\mathrm{s}}$, $\sigma_{8}]$ = [0.27[0.27[ 0.27, $0.05$, $0.67$, $0.96$, 0.83]0.83]0.83 ], evolves $N_{\mathrm{p}}=6912^{3}$ DM particles with mass resolution of $4.6\times 10^{8}\,h^{-1}M_{\odot}$ in a cubic box of width $L=1200\,h^{-1}\mathrm{Mpc}$. Both simulations assume flat cosmology, imposing $\Omega_{\Lambda}=1-\Omega_{\mathrm{m}}$, where $\Omega_{\mathrm{m}}=\Omega_{\mathrm{b}}+\Omega_{\mathrm{c}}$.
@@ 5.1 The shot noise effects on MFs
We find that the shapes of MFs are severely distorted by shot noises. The issue becomes increasingly conspicuous as the particle number density decreases. This is because particles successively become poorer tracers of the underlying DM field, making the downgraded fields choppier (cf. the top panels of Fig. 2). Low-density regions are more vulnerable to the down-sampling process, since shot noises mostly affect these regions. In particular, in poorly sampled regions, with sparse or even no particles, density fields are severely discontinuous or even blank (i.e., $01+\delta=$ 0) (cf. the top panels of Fig. 2). As particle number density decreases, blank regions become larger, thereby the volume fraction of non-zero density regions become smaller (cf. the left panel of Fig. 1 and top panels of Fig. 2). Due to the existences of blank regions, the regions with $01+\delta>$ 0 only account for a certain proportion of total box volume (cf. the left panel of Fig. 1). Thus, MFs have step changes [^7] Note that the values of MFs at $01+\delta=$ 0, which respectively are $V_{0}=1$, $V_{1}=0$, $V_{2}=0$, and $V_{3}=0$, cannot be plotted in Fig. 1, because of the logarithmic x-axis., when $01+\delta=0\Rightarrow 1+\delta>$ 0.
Moreover, shot noises give rise to various spurious structures on the excursion sets at different density thresholds, depending on down-sampling level. Specifically, for $10\%$ sampling, at low-density thresholds, more and larger isolated under-dense regions (bubbles) are produced by discrete-sampling process, such that $V_{1}$ and $V_{3}$ become larger and $V_{2}$ becomes smaller, relative to the case of noise-free field (the same below). At median density thresholds, we observe a lower negative minimum in $V_{3}$, indicating more porous structures (tunnels), produced by down sampling in the surfaces of excursion sets. As particle number density decreases, downgraded fields will be increasingly dominated by ‘meatball’ topology (i.e., a preponderance of isolated high-density regions). Because of this, for the extreme case of $0.1\%$ sampling, $V_{2}$ and $V_{3}$ are always positive when $11+\delta<$ 1. For the same case, we see that $V_{2}$ and $V_{3}$ have higher maximums at high-density thresholds, suggesting that the abundance of isolated regions is increased by breaking up structures into multiple objects as particles are taken out. For the intermediate case of $1\%$ sampling, the corresponding results naturally fall somewhere in between those of the former two cases.

### v2

@@ 1 Introduction
Ambitious on-going and up-coming cosmological surveys [e.g., HETDEX (Hill et al. 2008), LSST (LSST Science Collaboration et al. 2009), Euclid (Laureijs et al. 2011), 4MOST (de Jong et al. 2012), PFS (Takada et al. 2014), SPHEREX (Doré et al. 2014), DESI (DESI Collaboration et al. 2016), WFIRST (Doré et al. 2018), and CSST (Gong et al. 2019)], particularly the fifth-generation surveys [e.g., WST (Ellis & Dawson 2019), MSE (Percival et al. 2019), MegaMapper (Schlegel et al. 2019), and MUST [^1] MUltiplexed Survey Telescope (MUST): https://must.astro.tsinghua.edu.cn/must/en/index.html], will provide high-precision map of cosmic large-scale structure (LSS), which encodes a wealth of valuable cosmological information about our Universe. Efficient extractions of these critical information will help us greatly deepen our understanding of many key fundamental questions in cosmology (e.g., dark energy properties, neutrino masses, gravity and inflation models, etc.). This necessitates the development of powerful statistical tools to characterize or quantify LSS properties from various angles.
@@
Consequently, various summary statistics for non-Gaussian information have been proposed as potential supplements to two-point statistics, e.g., count in cells (de Lapparent et al. 1991; Uhlemann et al. 2020; Repp & Szapudi 2020), void statistics (Chan et al. 2014; Zhao et al. 2016; Kreisch et al. 2019), peak statistics (Gay et al. 2012; Shan et al. 2014; Liu et al. 2015a; Liu et al. 2015b), Voronoi statistics (Paranjape & Alam 2020; Jamieson & Loverde 2021), and even scattering transform (Cheng et al. 2020; Cheng & Ménard 2021), etc. In particular, morphological statistical methods [e.g., $\beta$-Skeleton (Fang et al. 2019; Suárez-Pérez et al. 2021), Betti numbers (Pranav et al. 2017; Pranav et al. 2019; Giri & Mellema 2021), persistent topology (Elbers & van de Weygaert 2019, 2023; Wilding et al. 2021; Jalali Kanafi et al. 2023; Bermejo et al. 2024; Yip et al. 2024), minimal spanning tree (Barrow et al. 1985; Naidoo et al. 2021), cosmic web skeleton (Novikov et al. 2006; Sousbie et al. 2008), wavelet analyses (Martinez et al. 1993; Arnalte-Mur et al. 2012), shape statistics (Sahni et al. 1998; Basilakos 2003), genus statistics (Gott et al. 1986; Hamilton et al. 1986; Appleby et al. 2021), Minkowski tensors (Appleby et al. 2018a; Appleby et al. 2018b), Minkowski functionals (Mecke et al. 1994; Liu et al. 2020; Matsubara et al. 2022), etc.] are actively employed to characterize the geometrical and topological properties of LSS, providing alternative ways to capture higher-order information directly, in complementary to traditional $N$-point formalism.
@@
As a conceptual generalization of genus (Mecke et al. 1994; Schmalzing & Buchert 1997), Minkowski functionals (MFs) can elegantly describe the global characterizations of morphological properties of LSS, and have been employed in various cosmological studies [e.g., detecting primordial non-Gaussianities (Hikage et al. 2006; Hikage et al. 2008; Planck Collaboration et al. 2016; Ducout et al. 2013), serving as standard ruler (Park & Kim 2010; Zunckel et al. 2011; Blake et al. 2014), constraining cosmological parameters (Appleby et al. 2018c; Appleby et al. 2021; Appleby et al. 2022), probing neutrino masses (Marques et al. 2019; Liu et al. 2020; Liu et al. 2022), analyzing effects of redshift space distortions (Codis et al. 2013; Jiang et al. 2021), testing cosmologies and gravities (Hikage et al. 2003; Wang et al. 2012; Fang et al. 2017), etc.]. One popular method [other method can also be found in literatures, i.e., the germ-grain approach (cf. Mecke et al. 1994; Wiegand et al. 2014)] for measuring MFs is the iso-density approach, which we will focus on in our study (see Section 2). The MFs measured in this way can be well theoretically modelled for Gaussian (Tomita 1990; Schmalzing & Buchert 1997; Matsubara 2003) and weakly non-Gaussian (Matsubara 1994; Matsubara 2003; Pogosyan et al. 2009; Matsubara 2010; Gay et al. 2012; Matsubara & Kuriki 2021; Matsubara et al. 2022) fields [even for the fields with RSDs (Matsubara 1996; Codis et al. 2013)], showing distinct advantages over the germ-grain approach.
@@
In this study, we opt for Delaunay Tessellation Field Estimator (DTFE) [^3] Delaunay Tessellation Field Estimator (DTFE): https://github.com/MariusCautun/DTFE (Schaap & van de Weygaert 2000; van de Weygaert & Schaap 2009; van de Weygaert et al. 2009; Cautun & van de Weygaert 2011a; Cautun & van de Weygaert 2011b) to address this issue [^4] Previously, Martínez et al. 2005 attempted to use wavelet-denoising method to alleviate the problem induced by fixed-kernel smoothing. Additionally, Zhang et al. 2010 endeavored to boost topological information of genus statistics with DTFE.. This method can help to recover the largest number of structural elements from point sets, allowing for extracting maximum amount of morphological information with MFs. Therefore, the statistical power of MFs with DTFE (hereafter DTFE MFs) can be significantly enhanced, over the fixed smoothing methods (Zhang et al. 2010). We first demonstrate this at matter field level and further extend it to halo field level. In particular, for DTFE MFs of halo fields, we numerically explore various main systematic effects, caused by finite voxel size, halo number density, halo-weighting scheme, and RSDs, on their measurements, and showcase their strong discriminative and constraining power in RSD studies. Hopefully, the applications of DTFE will dramatically improve the performance of MFs in extracting various key cosmological information (e.g., neutrino masses (Liu et al. 2020), modified gravities (Fang et al. 2017), primordial non-Gaussianities (Hikage et al. 2006; Hikage et al. 2008), and cosmological parameters (Appleby et al. 2020; Appleby et al. 2022), etc.) from sparse tracers.
@@ 2 Minkowski Functionals
MFs are a family of morphological (i.e., geometrical and topological) descriptors with properties of additivity, motion invariance, and conditional continuity for any manifold in $D$-dimensional space. These descriptors are originally derived from the theory of convex bodies and integral geometry, and were first introduced into cosmology by Ref. Mecke et al. 1994 to characterize the morphology (i.e., geometry and topology) of cosmic web. According to Hadwiger’s theorem (Hadwiger Springer, Berlin, 1957), the topo-geometrical properties of any given manifold in $D$-dimensional space can be completely characterized by $D+1$ MFs. Thus, it opens a unique way to comprehensively access all orders of N-point correlation information at once.
@@
Two standard grid-based numerical algorithms to compute MFs from regularly gridded density field [^5] In literatures, several novel triangulation-based algorithms (cf. Sheth et al. 2003; Aragon-Calvo et al. 2010; Lippich & Sánchez 2021), which estimate MFs by using triangulated iso-density surfaces, have also been proposed in succession, aiming to improve the accuracies of MFs’ estimations., i.e., Koenderink invariant from differential geometry (Schneider 1993) and Crofton’s formula from integral geometry (Hadwiger Springer, Berlin, 1957; Mecke 1979), have been developed in literatures (Schmalzing & Buchert 1997). In this work, we employ Crofton’s formula to quote our results, because it is more stable and is the most commonly used method in previous works. All MFs are measured as functions of density threshold $\nu\equiv 1+\delta$ with $\nu\in\left[0.003,1000\right]$, where $\delta$ is the density contrast. The error bars are estimated via the standard errors of MFs, i.e., $s_{\mathrm{e}}=\sigma/\sqrt{n_{\mathrm{f}}}$. Here, $\sigma$ is the standard deviation of MFs, calculated from $n_{\mathrm{f}}$ subfields obtained by equally subdividing the original field. For DM and halo fields, the numbers of subfields are $n_{\mathrm{f}}=4^{3}=64$ and $n_{\mathrm{f}}=8^{3}=512$, respectively. Since the error bars are too tiny to be visible in most scenarios, they are omitted in all figures in this paper, except for Fig. 12 and Fig. 13. Note that all MFs are visualized by logarithmic x-axis, considering that probability distribution function of LSS roughly obeys log-normal form at low redshift.
@@ 4 Cosmological $N$-body Simulations and Data Samples
In this work, we adopt two high-resolution pure cold dark matter (DM) $N$-body simulations for different application requirements. One simulation is dubbed as WMAP_$\_3072\_600$, which is from the CosmicGrowth simulation suite (Jing 2019), realized by running an adaptive parallel particle–particle–particle-mesh ($\mathrm{P}^{3}\mathrm{M}$) $N$-body code (Jing & Suto 2002; Jing et al. 2007). This simulation incorporates $N_{\mathrm{p}}=3072^{3}$ DM particles with mass resolution of $5.5\times 10^{8}\,h^{-1}M_{\odot}$ in a periodic cubic box with size of $L=600\,h^{-1}\mathrm{Mpc}$, adopting a WMAP cosmology, i.e., $[\Omega_{\mathrm{c}}$, $\Omega_{\mathrm{b}}$, $h$, $n_{\mathrm{s}}$, $\sigma_{8}]$ = [0.2235[0.2235[ 0.2235, $0.0445$, $0.71$, $0.968$, 0.83]0.83]0.83 ]. The other one is called TianZero simulation (Emberson et al. 2017; Yu et al. 2017), realized by using a publicly-available $\mathrm{P}^{3}\mathrm{M}$ $N$-body code CUBEP3M (Harnois-Déraps et al. 2013). This simulation, parameterized with $[\Omega_{\mathrm{c}}$, $\Omega_{\mathrm{b}}$, $h$, $n_{\mathrm{s}}$, $\sigma_{8}]$ = [0.27[0.27[ 0.27, $0.05$, $0.67$, $0.96$, 0.83]0.83]0.83 ], evolves $N_{\mathrm{p}}=6912^{3}$ DM particles with mass resolution of $4.6\times 10^{8}\,h^{-1}M_{\odot}$ in a cubic box of width $L=1200\,h^{-1}\mathrm{Mpc}$. Both simulations assume flat cosmology, imposing $\Omega_{\Lambda}=1-\Omega_{\mathrm{m}}$, where $\Omega_{\mathrm{m}}=\Omega_{\mathrm{b}}+\Omega_{\mathrm{c}}$.
@@ 5.1 The shot noise effects on MFs
We find that the shapes of MFs are severely distorted by shot noises. The issue becomes increasingly conspicuous as the particle number density decreases. This is because particles successively become poorer tracers of the underlying DM field, making the downgraded fields choppier (cf. the top panels of Fig. 2). Low-density regions are more vulnerable to the down-sampling process, since shot noises mostly affect these regions. In particular, in poorly sampled regions, with sparse or even no particles, density fields are severely discontinuous or even blank (i.e., $01+\delta=$ 0) (cf. the top panels of Fig. 2). As particle number density decreases, blank regions become larger, thereby the volume fraction of non-zero density regions becomes smaller (cf. the left panel of Fig. 1 and top panels of Fig. 2). Due to the existences of blank regions, the regions with $01+\delta>$ 0 only account for a certain proportion of total box volume (cf. the left panel of Fig. 1). Thus, MFs have step changes [^7] Note that the values of MFs at $01+\delta=$ 0, which respectively are $V_{0}=1$, $V_{1}=0$, $V_{2}=0$, and $V_{3}=0$, cannot be plotted in Fig. 1, because of the logarithmic x-axis., when $01+\delta=0\Rightarrow 1+\delta>$ 0.
Moreover, shot noises give rise to various spurious structures on the excursion sets at different density thresholds, depending on down-sampling level. Specifically, for $10\%$ sampling, at low-density thresholds, more and larger isolated under-dense regions (bubbles) are produced by discrete-sampling process, such that $V_{1}$ and $V_{3}$ become larger and $V_{2}$ becomes smaller, relative to the case of noise-free field (the same below). At median density thresholds, we observe a lower negative minimum in $V_{3}$, indicating more porous structures (tunnels), produced by down sampling in the surfaces of excursion sets. As particle number density decreases, downgraded fields will be increasingly dominated by “meatball topology” (i.e., a preponderance of isolated high-density regions). Because of this, for the extreme case of $0.1\%$ sampling, $V_{2}$ and $V_{3}$ are always positive when $11+\delta<$ 1. For the same case, we see that $V_{2}$ and $V_{3}$ have higher maximums at high-density thresholds, suggesting that the abundance of isolated regions is increased by breaking up structures into multiple objects as particles are taken out. For the intermediate case of $1\%$ sampling, the corresponding results naturally fall somewhere in between those of the former two cases.